### v2

# Communication-Efficient Soft Actor-Critic Policy Collaboration via Regulated Segment Mixture in Internet of Vehicles

Abstract
Multi-Agent Reinforcement Learning (MARL) has emerged as a foundational approach for addressing diverse, intelligent control tasks, notably in autonomous driving within the Internet of Vehicles (IoV) domain. However, the widely assumed existence of a central node for centralized, federated learning-assisted MARL might be impractical in highly dynamic environments. This can lead to excessive communication overhead, potentially overwhelming the IoV system. To address these challenges, we design a novel communication-efficient and policy collaboration algorithm for MARL under the frameworks of Soft Actor-Critic (SAC) and Decentralized Federated Learning (DFL), named RSM-MASAC, within a fully distributed architecture. In particular, RSM-MASAC enhances multi-agent collaboration and prioritizes higher communication efficiency in dynamic IoV system by incorporating the concept of segmented aggregation in DFL and augmenting multiple model replicas from received neighboring policy segments, which are subsequently employed as reconstructed referential policies for mixing. Distinctively diverging from traditional RL approaches, with derived new bounds under Maximum Entropy Reinforcement Learning (MERL), RSM-MASAC adopts a theory-guided mixture metric to regulate the selection of contributive referential policies to guarantee the soft policy improvement during communication-assisted mixing phase. Finally, the extensive simulations in mixed-autonomy traffic control scenarios verify the effectiveness and superiority of our algorithm.
Index Terms:
Communication-efficient, Multi-agent reinforcement learning, Soft actor critic, Regulated segment mixture, Internet of vehicles.
I Introduction
Internet of Vehicles (IoV) has evolved into a potent means to ubiquitously connect vehicles and enhance their self-driving capability. Various methods have been developed to control the speed of Connected Automated Vehicles (CAVs) in IoV to achieve better fleet management and accident avoidance, typically into three main types [2, 3, 4]: rule-based, optimization-based, and learning-based. Among these, learning-based Deep Reinforcement Learning (DRL) method has demonstrated superior driving performance [2, 5] and has become a widely discussed research area in IoV [6, 7, 8, 9, 10]. Specifically, CAVs can be modeled as DRL agents on top of a formulated Markov Decision Process (MDP). Correspondingly, these CAVs constitute a Multi-Agent Reinforcement Learning (MARL)-empowered system.
I-A Problem Statement and Motivation
In general, most MARL works adopt a Centralised Training and Decentralised Execution (CTDE) architecture [11, 12, 13, 14]. The model training process in IoV scenarios can be conducted either in the cloud or on a MEC server’s Vehicular Digital Twins (VDT) [15, 16] or through on-device Vehicular Federated Learning (VFL) [17, 18, 19]. Specifically, benefiting from Federated Learning (FL) by its continuous learning capabilities and periodic model updates (i.e. gradients or parameters) exchange, VFL maintains sensitive information, such as location, driving habits, vehicle status, and control policy models locally, reducing latency and enabling instantaneous decision-making. However, the dynamic challenges, including high mobility, intermittent connectivity, localized and decentralized nature of vehicles, make the common Centralized Federated Learning (CFL)’s assumption of a super central training controller impractical and underlies potential threat to the stability and timeliness of overall learning performance [20]. Besides, generalization difficulties [21, 22], such as variations in physical-environment (e.g., road changes or weather conditions) and social-environments (interactions with Human-Driven Vehicles (HDVs)), necessitate continuous training or fine-tuning of decision-making models for specific environments or tasks with minimal on-line data requirements when deployment.
In this way, there is growing research on peer-to-peer architecture’s Decentralized Federated Learning (DFL) [23] under VFL, focusing on the decentralized training/fine-tuning of Deep Neural Network (DNN) models among vehicles in real time [19, 24, 25]. This can be viewed as a distributed Stochastic Gradient Descent (SGD) optimization problem, in which the aggregation of exchanged DNN parameters acquired from periodic communication phase typically uses a simplistic parameter average mixture approach. In the context of Federated Reinforcement Learning (FRL) [26, 27, 28, 29], these exchanged and averaged parameters pertain to the RL agent’s policy network, which approximates the policy distribution through a DNN. And this can be implemented by Vehicle-to-Vehicle (V2V) communication as in most MARL works [30, 31, 32, 33]. However, the consequently frequent information exchange generates substantial communication overhead as the number of agents increases. Some DFL variants improve communication efficiency by reducing the number of communication rounds [34, 35, 36] or the communication workload per round [37, 24, 38, 39], but this can intensify the variability of local model updates, potentially leading to an inferior aggregated model post simple model parameter averaging [40]. This issue is more pronounced in online RL frameworks since RL agents interact more frequently with the environment than in offline supervised learning. The incremental arrival of data can magnify learning discrepancies among multiple agents and make vehicles less fault-tolerant. Interestingly, few studies have focused on the RL policies’ performance improvement after following this simple and straightforward aggregation of mixed policies using parameter averaging. However, not all communicated packets in MARL contribute effectively. During the communication-assisted mixing phase, MARL awaits a revolutionary policy parameter mixture method and corresponding metrics to regulate the aggregation of exchanged model updates, ensuring robust policy improvement for security and safety.
On the other hand, there is a contradiction between limited and costly sample availability and the need for fast learning speeds in driving policy, that is, MARL algorithms must fully explore and adapt to dynamic environments with minimal online learning data requirements. This aligns with the principles of Maximum Entropy Reinforcement Learning (MERL), such as Soft $Q$-learning [41] and Soft Actor Critic (SAC) [42, 43], which transforms the traditional reward maximum into both expected return and the expected entropy of the policy. The incorporated entropy in optimization redefines value function and policy optimization objective from the ground up. This not only enhances sample efficiency, adaptability, and robustness in the dynamic and uncertain IoV environment, but also overturns the proof of the traditional performance improvement bound [44, 45, 46, 47]. Therefore, this work is dedicated to reanalyzing efficient communication within the MERL and DFL framework. We propose a practical policy mixture method underpinned by a rigorously derived mixture metric, enabling MERL to achieve smooth and reliable parameter updates not only during independent local learning phase but also throughout the communication-assisted mixing phase.
Figure 1: Illustration of MARL in autonomous driving.
TABLE I: The Summary of Differences with Related Literature.
References Maximum Entropy Policy Improvement Guarantee Collaboration via Communication Efficient Communication Brief Description
[…]
[51, 52]
[…]
mately optimal RL’s performance guarantee analysis
[…]
I-B Contribution
In this paper, we propose the Regulated Segment Mixture-based Multi-Agent SAC (RSM-MASAC) algorithm, tailored for training local driving models of CAVs in the IoV using MERL while addressing communication overhead challenges inherent in DFL. Our primary contributions include:
  • For the highly dynamic IoV setting, the proposed RSM-MASAC algorithm effectively combines communication-efficient DFL with MERL. Within RSM-MASAC, vehicles capably receive segments of policy networks’ parameters from neighbors within the communication range only and constitute referential policies for selective parameter mixture. Due to the well-designed policy parameter mixture method, RSM-MASAC ensures credible performance improvement during the communication-assisted mixing phase while maintaining sufficient exploration in the local learning phase.
  • In terms of mixing a current policy and a referential policy, we derive a new, more generalized mixed policy improvement bound, which successfully tackles the analysis difficulties arising from the incorporation of redefined soft value function, dual policy optimization objective, and the logarithmic term of the policy for entropy maximum in MERL. Therefore, our work sets the stage for theoretically evaluating the performance of the mixed policy under distributed SGD optimization.
  • Instead of a simple parameter average in DFL, we bridge the relationship between DNN parameter gradient descent and MERL policy improvement, and regulate the selection of contributive referential policies by deriving a manageable, theory-guided mixture metric. Hence, it enhances the stability and practicality of directly mixing policy parameters during the communication-assisted mixing phase.
  • Through extensive simulations in the traffic speed control scenarios, our proposed RSM-MASAC algorithm could approach the converged performance of centralized FMARL [27] in a distributed manner, outperforming parameter average methods as in DFL [39, 24], thus confirming its effectiveness.
I-C Related Works
In Multi-Agent Systems (MAS), Independent Reinforcement Learning (IRL) [53, 54, 55], which relies solely on agents’ local perceptions without any collaboration, has been extensively studied and often used as baselines due to its varied policy performance, unstable learning and uncertain convergence [27]. Examples of IRL methods include IQL [54], IAC [13], IA2C [56] and IPPO [57], etc.
As an extension of IRL, distributed cooperative Reinforcement Learning (RL) enhances agents’ collective capabilities and efficiency by collaboratively seeking near-optimal solutions through limited information exchange with others. Notably, the exchanged contexts can be rather different and possibly include approximated value functions in [48, 49], rewards or even maximal $Q$-values on each state-action pair in [50]. Besides, given the clear evidence [58, 59] that experiences from homogeneous, independent learning agents in MAS can contribute to efficiently learning a commonly shared DNN model, the direct exchange of model updates during FL communication phase [26, 27, 60] is a viable way to indirectly integrate information and enhance cooperation among IRL. In addition to the commonly used centralized architecture that is less suitable for actual dynamic vehicular network, the peer-to-peer DFL paradigm, where clients exchange their local model updates only with their neighbors to achieve model consensus, emerges as an appealing alternative. This approach can be viewed as distributed SGD optimization, wherein the substantial communication expenditure cannot be overlooked. In that regard, some researchers resort to reduce communication frequency by aggregating more local updates before one round communication [34] or multi-round communication [35, 36]. Besides, either shrinking the number of parameters forwarded from local models or embarking on selective model synchronization is also tractable. For instance, Ref. [37] divides local gradients into several disjoint gradient partitions, with only a subset of these partitions being exchanged in any single communication round. Ref. [24] puts forward a randomized selection scheme for forwarding subsets of local model parameters to their one-hop neighbors. Furthermore, Ref. [38] suggests propagating top-k layers with higher normalized squared gradients, which may convey more information about the local data to neighbors. Meanwhile, Ref. [39] introduces a segmented gossip approach that involves synchronizing only model segments, thereby substantially splitting the communication expenditure.
Moreover, while the above distributed SGD optimization methods and the works in actual VFL [17, 18, 19] assume ideal communication between cooperative agents, ensuring exact and perfect information sharing without compromising privacy concerns may not be feasible with suboptimal wireless transmission links. Apart from complex, customized, and costly cryptographic schemes [40, 61], most works [62, 51, 52] suggest local quantization of data before transmission, which prevents the exposure of raw data. Additionally, this quantization, along with compression or sparsification [63, 36, 64], can significantly reduce the message size, while the corresponding convergence analyses have been detailed in [51, 35, 52].
Notably, when it comes to the method of mixing DNN parameters, these aforementioned DFL works generally adopt a simplistic averaging approach, which is familiar in parallel distributed SGD methods. However, when such an approach is directly applied into FRL, the crucial relationship between parameter gradient descent and policy improvement will be overshadowed [44]. Consequently, the corresponding mixture lacks proper, solid assessment means to prevent potential harm to policy performance [28]. In other words, directly using this kind of naive combination of communication efficient DFL and RL to enhance individual policy performance in IRL appears inefficient.
Regarding this issue, we can draw inspiration from the conservative policy iteration algorithm [45], which leverages the concept of policy advantage as a crucial indicator to gauge the improvement of the cumulative rewards, and applies a mixture update rule directly for policy distributions in its pursuit of an approximately optimal policy. Moreover, the mixture metric utilized in the update rule is also investigated to avoid making aggressive updates towards risky directions. This is particularly relevant as excessively large policy updates often lead to a significant deterioration of policy performance [28]. TRPO [47] substitutes the mixture metric with Kullback-Leibler (KL) divergence, a measure that quantifies the disparity between the distributions of the current and updated policies, and enables agents to learn monotonically improving policies. Ref. [46] extends this work into cooperative MARL settings. However, directly mixing policy distributions is often an intractable endeavor. To implement this procedure in practical settings with parameterised policies, Ref. [44] further simplifies the KL divergence to the parameter space through Fisher Information Matrix (FIM), so as to improve the policy performance by directly mixing DNN parameters. It focuses on stable policy updates throughout the communication-assisted mixing phase, but its analysis is confined to the traditional policy iteration-based RL algorithm, whose primary objective revolves around maximizing the expected return.
Transitioning to the realm of MERL, the integration of the entropy component marks a significant paradigm shift. MERL, exemplified by SAC, incorporates entropy maximization to encourage exploration. This dual objective of optimizing cumulative rewards and maintaining a high level of exploration fundamentally alters the criteria for policy improvement. As a result, traditional RL methodologies, particularly those based on PPO, are insufficient to address the new dynamics interplayed between reward maximization and exploratory behavior imposed by the entropy maximization in MERL. On the other hand, it’s also imperative to implement an appropriate and manageable mixture metric with monotonic policy improvement property to maximize the practicality of directly mixing policy parameters during the communication-assisted mixing phase. Such a metric must guarantee the efficacy of the resultant mixed policy in terms of policy improvement. Crucially, the assessment of the expected benefits of the mixed policy must be conducted before undertaking the actual policy mixing. That is, before any actual mixture of policy parameters, the potential of mixed policy to improve policy performance should be evaluated against the referential policy, whose parameters are received from neighbors in DFL. Only after confirming the anticipated benefits of the mixed policies should we proceed with the operation of mixing DNN parameters using the developed metric. Moreover, we have also summarized the key differences between our algorithm and the relevant literature in Table I. In conclusion, it is vital to reformulate the theoretical analysis within the combination of DFL and MERL framework, so as to offer robust theoretical underpinnings for practical applications in IoV.
I-D Paper Organization
The remainder of this paper is organized as follows. In Section II, we present preliminaries of SAC algorithm and main notations used in this paper. Then, we introduce the system model and formulate the problem in Section III. Afterwards, we provide the mixed performance improvement bound theorem of FRL communication under MERL in Section IV, and elaborate on the details of the proposed RSM-MASAC algorithm in Section V. In Section VI, we present the simulation settings and discuss the experimental results. Finally, Section VII concludes this paper.
II Preliminary
Beforehand, we summarize the mainly used notations in Table II.
TABLE II: Major notations used in the paper.
Notation Definition
$s_{t}^{(i)},a_{t}^{(i)},r^{(i)}_{t}$ Local state, individual action and reward of agent $i$ at time step $t$
$\pi,\tilde{\pi},\pi_{\text{mix}}$ Current target policy distribution, referential target policy distribution and the mixed policy distribution, which represents the probabilities of selecting each possible action given a state.
$\theta,\tilde{\theta},\theta_{\text{mix}}$ The parameter of specific neural network that parameterizes the target policy distribution, referential target policy distribution and the mixed policy distribution, respectively.
$p,P$ Index of segments, $p=1,2,\cdots,P$
$\Omega_{i}$ Set of one-hop neighbors within the communication range of agent $i$
$\zeta$ Mixture metric of current policy parameter vector and referential policy parameter vector
$\alpha$ Temperature parameter of policy entropy
$\varrho$ Target smoothing coefficient of target $Q$ networks
$\kappa$ Number of model replicas
$U$ Communication interval determined by specified iterations of the local policy
$\upsilon$ Transmission bits of the policy parameters
$\psi$   Communication consumption
We consider the standard RL setting, where the learning process of each agent can be formulated as an MDP. During the interaction with the environment, at each time step $t$, an RL agent observes a local state $s_{t}$ from local state space $\mathcal{S}$, and chooses an action $a_{t}$ from individual action space $\mathcal{A}$ according to the policy $\pi(\cdot|s_{t})\in\Pi$, which specifies a conditional distribution of all possible actions given the current state $s_{t}$, and $\Pi$ is the policy space. Then the agent receives an individual reward $r_{t}$, calculated by a reward function $\mathcal{R}:\mathcal{S}\times\mathcal{A}\rightarrow[0,R]$, and the environment transforms to a next state $s_{t+1}\sim p(\cdot|s_{t},a_{t})$. A trajectory starting from $s_{t}$ is denoted as $\tau_{t}=(s_{t},a_{t},s_{t+1},a_{t+1},\cdots)$. Besides, considering an infinite-horizon discounted MDP, the visitation probability of a certain state $s$ under the policy $\pi$ can be summarized as $d_{\pi}(s)=\sum\nolimits_{t=0}^{\infty}\gamma^{t}P(s_{t}=s;\pi)$, where $P(s_{t}=s;\pi)$ is the visitation probability of the state $s$ at time $t$ under policy $\pi$.
Different from traditional RL algorithms aiming to maximize the discounted expected total rewards only, SAC [42, 43] additionally seeks to enhance the expected policy entropy. To be specific, with the re-defined soft state value function $V^{\pi}(s_{0}):=\mathop{\mathbb{E}}\limits_{\tau_{0}\sim\pi}\left[\sum_{t=0}^{\infty}\gamma^{t}(r_{t}+\alpha H(\pi(\cdot|s_{t})))\right]$ on top of the policy entropy $H(\pi(\cdot|s))=-\mathbb{E}_{a\sim\pi}\log\pi(a|s)$, the task objective of SAC can be formulated as
$\vspace{-0.2em}\max\eta(\pi)=\mathop{\mathbb{E}}_{s_{0}\sim\rho_{0}}\left[V^{\pi}(s_{0})\right]$ (1)
where $\rho_{0}$ is the distribution of initial state $s_{0}$ and $\alpha\in(0,\infty)$ is a temperature parameter determining the relative importance of the entropy term versus the reward. Obviously, when $\alpha\rightarrow 0$, SAC gradually approaches the traditional RL. Meanwhile, the soft state-action value can be expressed as $Q^{\pi}(s_{t},a_{t}):=r_{t}+\mathop{\mathbb{E}}_{\tau_{t+1}\sim\pi}\!\left[\sum_{l=t+1}^{\infty}\gamma^{l-t}\Big{(}r_{l}\!+\!\alpha H(\pi(\cdot|s_{l}))\Big{)}\!\right]$ [42, 43], which does not include the policy entropy of current time step but includes the sum of all future policy entropy and the sum of all current and future rewards. Consistently, the state-action advantage value under policy $\pi$ is
$A_{\pi}(s_{t},a_{t})=Q^{\pi}(s_{t},a_{t})-V^{\pi}(s_{t})$ (2)
Accordingly, SAC maximizes (1) based on soft policy iteration, which alternates between soft policy evaluation and soft policy improvement.
  • For given $\pi$, soft policy evaluation implies that $Q^{\pi}$ is learned by repeatedly applying soft Bellman operator $\mathcal{T}^{\pi}$ to the real-valued estimate $Q$, given by:
    $\mathcal{T}^{\pi}Q(s_{t},a_{t})=r_{t}+\gamma\mathbb{E}_{s_{t+1}\sim p(\cdot|s_{t},a_{t})}[V(s_{t+1})],$ (3)
    where
    $V(s_{t})$ $=\mathbb{E}_{a_{t}\sim\pi}[Q(s_{t},a_{t})]+\alpha H(\pi(\cdot|s_{t}))$ (4)
    With $Q^{k+1}=\mathcal{T^{\pi}}Q^{k}$, as $k\rightarrow\infty$, $Q^{k}$ will converge to the soft $Q$ function $Q^{\pi}$ of $\pi$, as proven in [43].
  • In the soft policy improvement step, the goal is to find a policy $\pi_{\text{new}}$ superior than the current policy $\pi_{\text{old}}$, in terms of maximizing (1). Specifically, for each state, SAC updates the policy as
    $\pi_{\mathrm{new}}=$ $\mathop{\arg\min}_{\pi\in\Pi}\mathrm{D}_{\mathrm{KL}}\Big{(}\pi(\cdot|s_{t})\|\frac{\exp\big{(}\frac{1}{\alpha}Q^{\pi_{\text{old}}}({s_{t}},\cdot)\big{)}}{Z^{\pi_{\text{old}}}(s_{t})}\Big{)}$ (5)
    $=$ $\mathop{\arg\max}_{\pi\in\Pi}\mathbb{E}_{a_{t}\sim\pi}\left[Q^{\pi_{\text{old}}}(s_{t},a_{t})-\alpha\log\pi(a_{t}|s_{t})\right]$
    where $\mathrm{D}_{\mathrm{KL}}(\cdot)$ denotes the KL divergence, while the partition function $Z^{\pi_{\text{old}}}$ normalizes the distribution. Since $Z^{\pi_{\text{old}}}$ has no contribution to gradient with respect to the new policy, it can thus be ignored. As unveiled in Appendix A, through the update rule of (5), $Q^{\pi_{\text{new}}}(s_{t},a_{t})\geq Q^{\pi_{\text{old}}}(s_{t},a_{t})$ is guaranteed for all $(s_{t},a_{t})\in\mathcal{S}\times\mathcal{A}$. Notably, the proof of soft policy improvement detailed in Appendix A additionally serves as a confirmation for policy improvement of regulated segment mixture, which will be elaborated upon later.
Finally, with repeated application of soft policy evaluation and soft policy improvement, any policy $\pi\in\Pi$ will converge to the optimal policy $\pi^{*}$ such that $Q^{\pi^{*}}(s_{t},a_{t})\geq Q^{\pi}(s_{t},a_{t}),\forall(s_{t},a_{t})\in\mathcal{S}\times\mathcal{A}$, and the proof can be found in [42, 43].
III System Model and Problem Formulation
III-A System Model
As illustrated in Fig. 1, we primarily consider an IoV scenario consisting of $N$ CAVs, in which each CAV is controlled by an independent SAC-empowered agent, alongside some HDVs. Our MASAC learning encompasses an independent local learning phase and a communication-assisted mixing phase. In the first phase, we use SAC algorithm for each IRL agent $i$, $i\in\{1,2,\cdots,N\}$. That is, agent $i$ senses partial status $s_{t}^{(i)}$ (e.g., the speed and positions of neighboring vehicles) of the IoV environment, and has its local policy $\pi^{(i)}$ approximated by neural networks, and parameterized by $\theta\in\mathbb{R}^{d}$. It collects samples $\langle s_{t}^{(i)},a_{t}^{(i)},r_{t}^{(i)},s_{t+1}^{(i)}\rangle$ in replay buffer $\mathcal{D}^{(i)}$, and randomly samples a mini-batch $\Phi^{(i)}$ for local independent model updates [^1] Hereafter, for simplicity of representation, we omit the superscript $(i)$ under cases where the mentioned procedure applies for any agent.. Subsequently, in the second phase, each agent $i$ interacts with its one-hop neighbors $j\in\Omega_{i}$ within communication range, so as to reduce the behavioral localities of IRL and improve their cooperation efficiency.
III-A1 Local Learning Phase
Algorithmically, in the local learning phase with SAC, parameterized DNNs are used as approximators for policy and soft $Q$-function. Concretely, we alternate optimizing one policy network $\pi$ parameterized by $\theta$ and two soft $Q$ networks parameterized by $\omega_{1}$ and $\omega_{2}$, respectively. Besides, there are also two target soft $Q$ networks parameterized by $\bar{\omega}_{1}$ and $\bar{\omega}_{2}$, obtained as an exponentially moving average of current $Q$ network weights $\omega_{1}$, $\omega_{2}$. Yet, only the minimum $Q$ value of the two soft $Q$-functions is used for the SGD and policy gradient. This setting of two soft $Q$-functions will speed up training while the use of target $Q$ can stabilize the learning [42, 43, 65]. For training $\theta$ and $\omega\in{\omega_{1},\omega_{2}}$, agent randomly samples a batch of transition tuples from the replay buffer $\mathcal{D}$ and performs SGD. The parameters of each soft $Q$ network $\omega_{x},\forall x=1,2$ are updated through minimizing the soft Bellman residual error, that is,
$\vspace{-0.2em}J_{Q}(\omega_{x})=\frac{1}{2}\mathop{\mathbb{E}}\nolimits_{s_{t},a_{t}\sim\mathcal{D}}\left[Q_{\omega_{x}}({s}_{t},{a}_{t})-\hat{Q}(s_{t},a_{t})\right]^{2}\vspace{-0.2em}$ (6)
where the target $\hat{Q}(s_{t},a_{t})=r_{t}+\gamma\mathop{\mathbb{E}}_{\begin{subarray}{c}{s}_{t+1}\sim p(\cdot|s_{t},a_{t})\\ {a_{t+1}\sim\pi_{\theta}}\end{subarray}}\!\left[\mathop{\min}\limits_{x\in 1,2}Q_{\bar{\omega}_{x}}(s_{t+1},a_{t+1})\!-\!\alpha\log\pi_{\theta}(a_{t+1}|s_{t+1})\right]$.
Furthermore, the policy parameters of standard SAC can be learned according to (5) by replacing $Q^{\pi_{\text{old}}}$ with current $Q$ function estimate as
$J_{\pi}(\theta)=\!\mathop{\mathbb{E}}\limits_{{s}_{t}\sim\mathcal{D}}\!\big{[}\!\mathop{\mathbb{E}}\limits_{{a}_{t}\sim\pi_{\theta}}\!\big{[}\alpha\log\pi_{\theta}({a}_{t}|{s}_{t})\!-\!\mathop{\min}_{\begin{subarray}{c}x\in 1,2\end{subarray}}Q_{\omega_{x}}({s}_{t},{a}_{t})\big{]}\big{]}\vspace{-0.2em}$ (7)
Besides, since the gradient estimation of (7) has to depend on the actions stochastically sampled from $\pi_{\theta}$, which leads to high gradient variance, the reparameterization trick [42, 43] is used to transform the action generation process into deterministic computation, allowing for efficient gradient-based training. Concretely, the random action $a_{t}$ can be expressed as a reparameterized variable
$\vspace{-0.2em}a_{t}=f_{\theta}(s_{t};\delta_{t})=\tanh\big{(}\mu_{\theta}(s_{t})+\delta_{t}\odot\sigma_{\theta}(s_{t})\big{)}\vspace{-0.2em}$ (8)
where $\odot$ represents Hadamard product, $\delta_{t}$ is sampled from $\mathcal{N}(0,\mathbf{I}_{\dim\mathcal{A}})$, the mean $\mu_{\theta}$ and standard $\sigma_{\theta}$ are outputs from the policy network parameterized by $\theta\in\mathbb{R}^{d}$. This reparameterization enables the action $a_{t}$ to be a differentiable function of $\theta$, facilitating gradient descent methods.
In addition to the soft $Q$-function and the policy, the temperature parameter $\alpha$ can be automatically updated by optimizing the following loss
$J(\alpha)=\mathop{\mathbb{E}}\limits_{s_{t}\sim\mathcal{D}}\big{[}\mathop{\mathbb{E}}_{{a}_{t}\sim\pi_{\theta}}[-\alpha\log\pi_{\theta}({a}_{t}|{s}_{t})-\alpha\bar{\mathcal{H}}]\big{]}\vspace{-0.2em}$ (9)
where $\bar{\mathcal{H}}$ is an entropy target with default value $-\dim{\mathcal{A}}$. Thus the policy can explore more in regions where the optimal action is uncertain, and remain more deterministic in states with a clear distinction between good and bad actions.
In addition, since two-timescale updates, i.e, less frequent policy updates, usually result in higher quality policy updates. Therefore, we integrate the delayed policy update mechanism employed in TD3 [66], into our framework’s methodology. To this end, the policy, temperature and target $Q$ networks are updated with respect to soft $Q$ network every $e$ iterations.
III-A2 Communication-Assisted Mixing Phase
Subsequent to the phase of local learning, neighboring agents initiate periodic communication to enhance their collaboration in accomplishing complex tasks. The messages transmitted by agents are limited to policy parameters $\theta$ as in VFL [17, 18, 19]. Such an assumption is feasible as soft $Q$-functions have less impact on action selection than the policy in actor-critic algorithm.
Specifically, every $U$ times of policy updates, a communication round begins. Agent $i$ receives the policy parameters $\theta^{(j)}$ from neighboring agents $j\in\Omega_{i}$ via the V2V collaboration channel. Subsequently, agent $i$ could employ various methods to formulate a referential policy $\tilde{\pi}^{(i)}$, which is parameterized by $\tilde{\theta}^{(i)}\!=\!f({\theta}^{(1)},\cdots,{\theta}^{(j)},\cdots)$, that is, the parameters obtained from its neighbors $\forall j\in\Omega_{i}$.
Afterwards, agent $i$ directly mixes DNN parameters $\theta$ of the policy network, consistently with parallel distributed SGD methods as
$\theta_{\text{mix}}^{(i)}=\theta^{(i)}+\zeta(\tilde{\theta}^{(i)}-\theta^{(i)})\vspace{-0.1em}$ (10)
where $\zeta\in[0,1]$ is the mixture metric of DNN parameters. Taking model averaging in [35, 44] as example, $\tilde{\theta}^{(i)}$ is computed as $\tilde{\theta}^{(i)}\!=\!\sum\nolimits_{j\in\Omega_{i}}\!{\theta}^{(j)}$, and $\zeta\!=\!1\!-\!1/|\Omega_{i}|$ is further influenced by the number of neighbors involved. Then, for each agent $i$, $\theta^{(i)}$ should get aligned with mixed policy’s parameters $\theta_{\text{mix}}^{(i)}$.
III-B Problem Formulation
This paper primarily targets the communication assisted mixing phase in MASAC. Instead of simply continuing to follow the idea of direct average of the DNN parameters in FL, an effective parameter mixture method could better leverage the exchanged parameters to yield a superior target referential policy. This approach aims to improve the reinforcement learning policy performance during communication assisted mixing phase, resulting in a consistently higher value with respect to the maximum entropy objective, as described in (1). However, it remains little investigated on the feasible means to mix the exchanged parameters (or their partial segments) and determine the proper mixture metric in (10), though it vitally affects both the communication overhead and learning performance. Therefore, by optimizing the mixture metric $\zeta$, we mainly focus on reducing the communication expenditure while maintain acceptable cumulative rewards, that is,
$\min_{\zeta}c(\upsilon,f)$
$s.t.\quad$ $\sum\nolimits_{t}r_{t}(\Theta,\zeta)\geq r^{\text{thre}}$
$\Theta\leftarrow\{\theta_{\text{mix},k}^{(1)},\cdots,\theta_{\text{mix},k}^{(N)}\}$ (11)
$\theta_{\text{mix},k}^{(i)}=\theta_{k}^{(i)}+\zeta(\tilde{\theta}_{k}^{(i)}-\theta_{k}^{(i)}),$ $\forall i\in\{1,\cdots,N\}$
$\tilde{\theta}_{k}^{(i)}=f({\theta}_{k}^{(1)},\cdots,{\theta}_{k}^{(j)},\cdots),$ $\forall k\text{ mod }U=0,j\in\Omega_{i}$
where $r^{\text{thre}}$ denotes the required minimum cumulative rewards, $v$ indicates the transmission bits of policy parameters, and $k$ is the index of policy iterations. Furthermore, $c(v,f)$ denotes the communication expenditure, governed by the utilization function $f$, which represents the method of utilizing the parameter or parameter segments obtained from different neighboring agent $j\in\Omega_{i}$ to construct a referential policy parameterized by $\tilde{\theta}$. The practical implementation of $f$ can be contingent on various factors, i.e., the underlying communications capability of vehicles to simultaneously receive and handle signals and the environmental conditions. After obtaining the referential policy parameter $\tilde{\theta}$, it is worthwhile to resort to a more comprehensive design of $\zeta$ to calibrate the communicating agents and contents as well as regulate the means to mix parameters or parameter segments, so as to provide a guarantee of performance improvement.
IV Mixed Performance Improvement Bound Theorem of FRL Communication under MERL
(a) Independent local learning phase
(b) Communication-assisted mixing phase
Figure 2: The illustration of RSM-MASAC implementation.
Given the potential for significant variability in policy performance due to differences in training samples among multiple agents, as well as the staleness of iterations caused by varying computing power of different vehicles, it is critical to recognize that not all referential policies - those amalgamated with the agent’s own policy via the DNN parameter mixture approach detailed in (10) - can contribute positively to agent’s local learning process. More seriously, it may even degrade the learning performance sometimes [44, 46]. Therefore, in order to ensure the policy improvement post-mixture, a robust theoretical analysis method for evaluating the efficacy of the mixed policy distribution $\pi_{\text{mix}}$, approximated by DNN with parameter $\theta_{\text{mix}}$, is essential. This would regulate the selection of only those referential policies $\tilde{\pi}$, parameterized by $\tilde{\theta}$, that are beneficial, into the parameter mixture process.
Drawing from this premise, based upon the conservative policy iteration as outlined in [45], we employ a mixture update rule on policy distributions to find an approximately optimal policy. Our analysis here will focus specifically on calibrating the mixing means of policy distributions $\pi$ that exhibit a monotonic policy improvement property during the communication-assisted mixing phase. Notably, due to the non-linear transformation in DNN and possible applicability of softmax or restricted reparameterization [67], the mixture of policy distribution in (12) is not directly equivalent to the mixture of their parameters, especially when the parameter mixture takes the form in (10) under distributed SGD. Hence, the detailed and more practical mixture implementation for the policy network parameter $\theta$, as in (10), will be derived from this section and more thoroughly presented in Section V-A2.
For any state $s$, we also define the mixed policy $\pi_{\text{mix}}$, which refers to a mixed distribution, as the linear combination of any referential policy $\tilde{\pi}$ and current policy $\pi$
$\pi_{\text{mix}}(a|s)=(1-\beta)\pi(a|s)+\beta\tilde{\pi}(a|s)$ (12)
where $\beta\in[0,1]$ is the weighting factor. The soft policy improvement, as outlined in Appendix A, suggests that this policy mixture can influence the ultimate performance regarding the objective stated in (1). Fortunately, we have the following new theorem on the performance gap associated with adopting these two different policies.
Theorem 1.
(Mixed Policy Improvement Bound) For any policy $\pi$ and $\tilde{\pi}$ adhering to (12), the improvement in policy performance after mixing can be measured by:
$\eta(\pi_{\text{mix}})-\eta(\pi)$ $\geq\beta\mathop{\mathbb{E}}_{\begin{subarray}{c}s\sim d_{\pi}\\ a\sim\tilde{\pi}\end{subarray}}\left[A_{\pi}(s,a)+\alpha H(\tilde{\pi}(\cdot|s))\right]-\frac{2\gamma\varepsilon\beta^{2}}{(1-\gamma)^{2}}$
$\quad+\alpha\mathop{\mathbb{E}}_{s\sim d_{\pi_{\text{mix}}}}\left[\mathrm{D}_{\mathrm{JS}}^{\beta}(\tilde{\pi}(\cdot|s)\|\pi(\cdot|s))\right]$ (13)
where $\varepsilon\!:=\!\max_{s}|\mathop{\mathbb{E}}_{a\sim\tilde{\pi}}[A_{\pi}(s,a)\!+\!\alpha H(\tilde{\pi}(\cdot|s))]|$ represents the maximum advantage of $\tilde{\pi}$ relative to $\pi$, and $\mathrm{D}_{\mathrm{JS}}^{\beta}(p\|q)=\beta\sum p\log\frac{p}{\beta p+(1-\beta)q}+(1-\beta)\sum q\log\frac{q}{\beta p+(1-\beta)q}$ is the $\beta$-skew Jensen-Shannon(JS)-symmetrization of KL divergence [68], with two distributions $p$ and $q$.
The proof of this theorem is given in Appendix B.
Remark: Notably, the proof effectively tackles the difficulties arising from the re-defined, soft state value function with the extra logarithmic term of the policy by utilizing entropy decomposition of $H(\pi_{\text{mix}}(\cdot|s))$ in Lemma 5 as well as several mathematical tricks. Hence, this sets the stage for theoretically evaluating the performance of the mixed policy before the mixture occurs. The mixed policy improvement bound in (1) implies that under the condition that the right-hand side of (1) is larger than zero, the mixed policy will assuredly lead to an improvement in the true expected objective $\eta$. Besides, from another point of view, any mixed policy with a guaranteed policy improvement in Theorem 1 definitely satisfies the soft policy improvement as well, since the item $\mathbb{E}_{s_{1}}\left(\alpha H(\pi_{\text{old}}(\cdot|s_{1}))\!+\!\mathop{\mathbb{E}}_{\begin{subarray}{c}a_{1}\sim\pi_{\text{old}}\end{subarray}}[Q^{\pi_{\text{old}}}(s_{1},a_{1})]\right)=\mathbb{E}_{s_{1}}[V^{\pi_{\text{old}}}(s_{1})]$ in (21) in the Appendix A (i.e., proof of Lemma 1) is less than $\mathbb{E}_{s_{1}}[V^{\pi_{\text{mix}}}(s_{1})]$. In addition, when the temperature parameter $\alpha=0$, this inequality can reduce to the standard form of policy improvement in traditional RL [45, 47, 44]. Hence, Theorem 1 derives a more general conclusion for both MERL and traditional RL.
Furthermore, since for $\beta\in[0,1]$, $\mathrm{D}_{\text{JS}}^{\beta}$ is greater than zero [68], only the sign of the first two terms in (1) need to be considered. By applying the re-defined policy advantage under MERL
$\mathbb{A}^{+}_{\pi}(\tilde{\pi}):=\mathop{\mathbb{E}}_{\begin{subarray}{c}s\sim d_{\pi},a\sim\tilde{\pi}\end{subarray}}\left[A_{\pi}(s,a)+\alpha H(\tilde{\pi}(\cdot|s))\right]$ (14)
we can therefore establish a tighter yet more tractable bound for policy improvement as
$\eta(\pi_{\text{mix}})-\eta(\pi)$ $\geq\beta\mathbb{A}^{+}_{\pi}(\tilde{\pi})-C\beta^{2}$ (15)
where $C=\frac{2\varepsilon\gamma}{(1-\gamma)^{2}}$. Thus, (15) indicates that a mixed policy conforms to the principle of soft policy improvement, provided that the right side of (15) yields a positive value. More specifically, if the policy advantage $\mathbb{A}^{+}_{\pi}(\tilde{\pi})$ is positive, an agent with policy $\pi$ can reap benefits by mixing its policy distribution with referential policy $\tilde{\pi}$. On the contrary, if this advantage is non-positive, policy improvement cannot be assured through the mixing of $\tilde{\pi}$ and $\pi$. In essence, (15) serves as a criterion for selecting referential policies that ensure final performance improvement, and thus we have the following corollary.
Corollary 1.
To obtain guaranteed performance improvement, the mixture approach of policy distributions shall satisfy that
  • $\mathbb{A}^{+}_{\pi}(\tilde{\pi})>0$;
  • $\beta\mathbb{A}^{+}_{\pi}(\tilde{\pi})-C\beta^{2}>0$.
V MASAC with Regulated Segment Mixture
In this section, as shown in Fig. 2, we present the design of RSM-MASAC, which reduces the communication overhead while incurring little sacrifice to the learning performance.
V-A Algorithm Design
Consistent with the SAC setting as in Section III, agents in RSM-MASAC undergo the same local iteration process. Meanwhile, for the communication-assisted mixing phase, we will elaborate on the design details of RSM-MASAC, including the segment request & response and policy parameter mixture with theory-established performance improvement.
V-A1 Segment Request & Response
Inspired by segmented pulling synchronization mechanism in DFL [69], we develop and perform a segment request & response procedure. The proposed approach divides transmission of the policy parameters into segments, and each agent selectively request various segments of policy parameters from different neighbors simultaneously through V2V communication, thereby facilitating the construction of a reconstructed referential policy for subsequent aggregation while also effectively balancing the load of communication costs and optimizing bandwidth usage. Specifically, for every communication round, each agent $i$ breaks its policy parameters $\theta^{(i)}$ into $P$ ($P\leq|\Omega_{i}|$, according to the number of different agents’ neighbors in current time) non-overlapping segments $\hat{\theta}^{(i)}_{1},\hat{\theta}^{(i)}_{2},\cdots,\hat{\theta}^{(i)}_{P}$ as
$\theta^{(i)}=(\hat{\theta}^{(i)}_{1},\hat{\theta}^{(i)}_{2},\cdots,\hat{\theta}^{(i)}_{P})\vspace{-0.1em}$ (16)
Significantly, the available segmentation strategies are diverse and include, but are not limited to, dividing the policy parameters by splitting DNN layers [38], modular approach [70], or segmenting according to the parameter size [69]. Meanwhile, there exist rather diverse V2V communication protocols and technologies such as DSRC (based on IEEE 802.11p and its subsequent IEEE 802.11bd), NR-V2X (as specified in 3GPP Release 16 and 17), which support Multiple Input Multiple Output (MIMO) technology and advanced signal processing techniques, such as spatial multiplexing and beamforming. Therefore, the communication capability depends on multiple factors ranging from physical layer configurations, such as antenna array setup and channel conditions, to upper layer protocols, shaping the theoretical upper limit for segmentation granularity $P_{\max}$. For instance, a $4\times 4$ MIMO system in a vehicle can theoretically support up to $4$ independent data streams of policy segments simultaneously, enhancing data throughput and robustness. Besides, shorter distances between vehicles and better channel conditions generally enhance the capacity and reliability of system. Still, the segmentation also can be dynamic, with larger $P_{\max}$ in environments with higher vehicle density and smaller $P_{\max}$ in cases where fewer vehicles are within communication range.
Here, to clarify this process, we use the most intuitive uniform parameter partition. For each segment $p=1,\cdots,P$, agent $i$ randomly selects a target agent (without replacement) from its neighbors (i.e., $j_{p}\in\Omega_{i}$) to send segment request $(i,P,j_{p},p)$, which indicates the agent $i$ who initiates the request and its total segment number $P$, as well as the requested segment $p$ from the target agent $j_{p}$. Upon receiving the request, the agent $j_{p}$ will break its own policy parameters $\theta^{(j_{p})}$ into $P$ segments and return the corresponding requested segment $\hat{\theta}^{(j_{p})}_{p}$ according to the identifier $p$. The number of segments $P$ can be flexibly determined according to the actual communication range of agent $i$, i.e., $P=\min\{P,|\Omega_{i}|\}<P_{\max}$. It should be noted that in order to reduce the complexity and facilitate the implementation, we only discuss the case as in Fig. 2 that $P$ is the same constant for all agents and is not greater than $\max_{i}|\Omega_{i}|,\forall i$. Meanwhile, the segmentation methodology is identical across agents. Then, agent $i$ could reconstruct a referential policy based on all of the fetched segments, that is,
$\tilde{\theta}^{(i)}=(\hat{\theta}^{(j_{1})}_{1},\hat{\theta}_{2}^{(j_{2})},\cdots,\hat{\theta}_{P}^{(j_{P})})\vspace{-0.1cm}$ (17)
In fact, this segmented transmission approach can be executed in parallel, thereby optimizing the utilization of available bandwidth. Instead of being confined to a single link, the traffic is distributed across $P$ links, enhancing the overall data transfer efficiency. Besides, in order to further accelerate the propagation and ensure the model quality, we can construct multiple model replicas in RSM-MASAC. That is, the process of segment request &\&& response can be repeated $P\times\kappa$ times, reconstructing $\kappa$ reconstructed referential policies in one communication round.
Algorithm 1 RSM-MASAC Algorithm
1:  Initialize network parameters $\theta^{(i)}$, $\omega_{1}^{(i)}$, $\omega_{2}^{(i)}$, $i=1,2,\cdots N$
2:  Initialize target network parameters $\bar{\omega}_{1}^{(i)}\leftarrow\omega_{1}^{(i)}$, $\bar{\omega}_{2}^{(i)}\leftarrow\omega_{2}^{(i)}$
3:  Initialize learning rate $\eta_{\pi}$, $\eta_{Q}$, $\eta_{\alpha}$, temperature $\alpha$, communication interval $U$, number of segments $P$ and number of replicas $\kappa$
4:  Initialize iteration index $k\leftarrow 0$ and $counter\leftarrow 0$
5:  for each epoch do
6:     for $t\leftarrow 1$ to $T$ do
7:        Each agent $i$ executes:
8:        * Independent local learning phase
9:        Select an action $a_{t}^{(i)}$ with respect to $s_{t}^{(i)}$ according to the current policy $\pi^{(i)}$ parameterized by $\theta^{(i)}$
10:        Observe reward $r_{t}^{(i)}$ and next state $s_{t+1}^{(i)}$
11:        Save the new transition in replay buffer: $\mathcal{D}^{(i)}\leftarrow\mathcal{D}^{(i)}\cup\langle s_{t}^{(i)},a_{t}^{(i)},r_{t}^{(i)},s_{t+1}^{(i)}\rangle$
12:        Sample a mini-batch $\Phi^{(i)}\sim\mathcal{D}^{(i)}$
13:        Update soft $Q$ function $\omega_{x}^{(i)}\leftarrow\omega_{x}^{(i)}-\eta_{Q}\nabla J_{Q}(\omega_{x}^{(i)})$ by (6), for $\forall x={1,2}$
14:        if $counter\bmod d=0$ or $t=T$ then
15:           Update policy $\vspace{-0.5em}\theta_{k+1}^{(i)}\leftarrow\theta_{k}^{(i)}-\eta_{\pi}\nabla J_{\pi}(\theta_{k}^{(i)})$ by (7)
16:           Adjust temperature $\alpha^{(i)}\leftarrow\alpha^{(i)}-\eta_{\alpha}\nabla J(\alpha^{(i)})$ by (9)
17:           Update target networks $\bar{\omega}_{x}^{(i)}\leftarrow\varrho\omega_{x}^{(i)}+(1-\varrho)\bar{\omega}_{x}^{(i)}$, for $\forall x=1,2$
18:           $k\leftarrow k+1$
19:        end if
20:        $counter\leftarrow counter+1$
21:        * Communication-assisted mixing phase
22:        if $k\bmod U=0$ then
23:           Update policy $\theta_{k}^{(i)}\leftarrow$ CommMix($\theta_{k}^{(i)}$, $i$, $\Omega_{i}$, $P$, $\kappa$)
24:        end if
25:     end for
26:  end for
27:  CommMix($\theta,i,\Omega_{i},P,\kappa$):
28:       for each replica $u=1,2,\cdots,\kappa$ do
29:           Send $P$ pulling request $(i,P,j_{p},p)$ to nearby collaborators   in $\Omega_{i}$, and receive $\hat{\theta}^{(j_{p})}_{p}$ to reconstruct $\tilde{\theta}$ as (17)
30:           Select $M$ samples from the replay buffer $\mathcal{D}^{(i)}$
31:           Estimate $\mathbb{A}^{+}_{\pi}(\tilde{\pi})$ according to (19)
32:           if $\mathbb{A}^{+}_{\pi}(\tilde{\pi})>0$ then
33:               Evaluate $F(\theta)$ according to (20)
34:               Get the upper bound of $\zeta$ according to Theorem 2
35:               Make the mixture metric $\zeta$ less than the calculated upper    bound, and update $\theta_{\text{mix}}^{(i)}$ by (10)
36:              $\quad\theta\leftarrow\theta_{\text{mix}}^{(i)}$
37:           end if
38:       end for
V-A2 Policy Parameter Mixture with Theory-Established Performance Improvement
Consistent with TRPO [47], we introduce KL divergence to replace $\beta$ by setting $\beta:=\sqrt{\mathrm{D}_{\mathrm{KL}}^{\max}(\pi\|\pi_{\text{mix}})}$, where $\mathrm{D}_{\mathrm{KL}}^{\max}(\pi\|\pi_{\text{mix}})=\max_{s}\mathrm{D}_{\mathrm{KL}}(\pi(\cdot|s)\|\pi_{\text{mix}}(\cdot|s))$. Thus, the second condition in Corollary 1 is equivalent to
$\vspace{-0.2em}\sqrt{\mathrm{D}_{\mathrm{KL}}^{\max}(\pi\|\pi_{\text{mix}})}<\frac{\mathbb{A}^{+}_{\pi}(\tilde{\pi})}{C}.\vspace{-0.2em}$ (18)
Nevertheless, it hinges on the computation-costly KL divergence to quantify the difference between probability distributions. Fortunately, since for a small change in the policy parameters, the KL divergence between the original policy and the updated policy can be approximated using a second-order Taylor expansion, wherein FIM serves as the coefficient matrix for the quadratic term. This provides a tractable way to assess the impact of parameter changes on the policy. Therefore, we utilize FIM, delineated in context of natural policy gradients by [71], as a mapping mechanism to revise the impact of certain changes in policy parameter space on probability distribution space. Then in the following theorem, we can get the easier-to-follow, trustable upper bound for the mixture metric of policy DNN parameters.
Theorem 2.
(Guaranteed Policy Improvement via Parameter Mixing) With any referential policy parameters $\tilde{\theta}$, an agent with current policy parameters $\theta$ can improve the true objective $\eta$ as in (1) through updating $\theta$ to mixed policy parameters $\theta_{\text{mix}}$ in accordance with (10), provided it fulfills the following two conditions:
  • $\mathbb{A}^{+}_{\pi}(\tilde{\pi})>0$;
  • $0<\zeta<\left[\frac{2\mathbb{A}^{+}_{\pi}(\tilde{\pi})}{C[(\tilde{\theta}-\theta)^{\intercal}F(\theta)(\tilde{\theta}-\theta)]}\right]^{\frac{1}{2}}$.
where $F(\theta)$ is the FIM of policy $\pi$ parameterized by $\theta$.
Proof.
Recalling the definition of $\beta$ and the mixture approach of $\theta_{\text{mix}}$ in (10), as well the change of policy parameters (i.e., $\Delta\theta=\zeta(\tilde{\theta}-\theta)$), for any state $s$, the KL divergence between the current policy and the mixed policy can be expressed by performing a second-order Taylor expansion of the KL divergence at the point $\theta$ in parameter space as
$\vspace{-3em}\mathrm{D}_{\mathrm{KL}}$ $(\pi_{\theta}(\cdot|s)\|\pi_{\theta+\Delta\theta}(\cdot|s))$
$=$ $\sum_{a\in\mathcal{A}}\pi_{\theta}(a|s)\log\frac{\pi_{\theta}(a|s)}{\pi_{\theta+\Delta\theta}(a|s)}$
$\approx$ $\mathrm{D}_{\mathrm{KL}}(\pi_{\theta}(\cdot|s)\|\pi_{\theta}(\cdot|s))-\mathop{\mathbb{E}}\limits_{a\sim\pi_{\theta}}\left[\frac{\partial\log\pi_{\theta}(a|s)}{\partial\theta}\right]^{\intercal}\!\!\Delta\theta$
$\quad+\frac{1}{2}\Delta\theta^{\intercal}F(\theta)\Delta\theta$
=$\stackrel{{\scriptstyle(a)}}{{=}}$ $\frac{1}{2}\zeta^{2}(\tilde{\theta}-\theta)^{\intercal}F(\theta)(\tilde{\theta}-\theta)\vspace{-2em}$
where the equality $(a)$ comes from the fact that by definition, $\mathrm{D}_{\mathrm{KL}}(\pi_{\theta}(\cdot|s)\|\pi_{\theta}(\cdot|s))=0$, while $\mathop{\mathbb{E}}\nolimits_{a\sim\pi_{\theta}}\left[\frac{\partial\log\pi_{\theta}(a|s)}{\partial\theta}\right]=\sum_{a\in\mathcal{A}}\pi_{\theta}(a|s)\frac{\partial\log\pi_{\theta}(a|s)}{\partial\theta}=\sum_{a\in\mathcal{A}}\frac{\partial\pi_{\theta}(a|s)}{\partial\theta}=0$. Notably, the FIM $F(\theta)$ takes the expectation for all possible states, which reflects the average sensitivity of the whole state space rather than a particular state, and can be calculated as
$\vspace{-0.5em}F(\theta)=\mathop{\mathbb{E}}_{\begin{subarray}{c}s\sim d_{\pi}\\ a\sim\pi_{\theta}\end{subarray}}\left[\left(\frac{\partial\log\pi_{\theta}(a|s)}{\partial\theta}\right)\!\left(\frac{\partial\log\pi_{\theta}(a|s)}{\partial\theta}\right)^{\intercal}\right]$
Finally, combining Corollary 1 and (18), we have the theorem. ∎
Remark: With Theorem 1 and Theorem 2, we can anticipate the extent of policy performance changes resulting from policy DNN parameters mixture during the communication-assisted mixing phase. This prediction is based solely on the local curvature of the policy space provided by FIM, thus avoiding the cumbersome computations to derive the full KL divergence of two distributions for every state. Besides, given that $F(\theta)$ is a positive definite matrix, a positive sign of the policy advantage $\mathbb{A}^{+}_{\pi}(\tilde{\pi})$ implies an increase in the DNN parameters’ mixture metric $\zeta$ corresponding to the rise in $\mathbb{A}^{+}_{\pi}(\tilde{\pi})$. That is, rather than simple averaging, agents can achieve guaranteed soft policy improvement after mixing by learning more effectively from referential policies with an elevated $\zeta$.
In practice, to evaluate $\mathbb{A}^{+}_{\pi}(\tilde{\pi})$ and $F(\theta)$, the expectation can be estimated by the Monte Carlo method, approximating the global average by the states and actions under the policy. Meanwhile, the importance sampling estimator is also adopted to use the off-policy data in the replay buffer for the policy advantage estimation, where $\pi_{t}$ typically denotes the action sampling policy at time step $t$. As outlined in Lemma 7 in Appendix C, we have
$\mathbb{A}^{+}_{\pi}(\tilde{\pi})$ $\approx\mathbb{E}_{s_{t},a_{t}\sim\mathcal{D}}\bigg{[}\left(\frac{\tilde{\pi}_{\tilde{\theta}}(a_{t}|s_{t})-\pi_{\theta}(a_{t}|s_{t})}{\pi_{t}(a_{t}|s_{t})}\right)\!\mathop{\min}_{\begin{subarray}{c}x\in 1,2\end{subarray}}Q_{\omega_{x}}({s}_{t},{a}_{t})$
$\qquad+\alpha[H(\tilde{\pi}_{\tilde{\theta}}(\cdot|s_{t}))-H({\pi}_{{\theta}}(\cdot|s_{t}))]\bigg{]}$ (19)
and
$F(\theta)\!\approx\!\mathop{\mathbb{E}}\limits_{s_{t}\sim\mathcal{D}}\!\bigg{[}\!\mathop{\mathbb{E}}\limits_{a_{t}\sim\pi_{\theta}}\!\bigg{[}\!\left(\frac{\partial\log\pi_{\theta}(a|s)}{\partial\theta}\right)\!\left(\frac{\partial\log\pi_{\theta}(a|s)}{\partial\theta}\right)^{\intercal}\!\bigg{]}\bigg{]}$ (20)
Finally, we summarize the details of RSM-MASAC in Algorithm 1. RSM-MASAC employs a theory-guided metric for policy parameter mixture, which takes into account the potential influence of parameters mixture on soft policy improvement under MERL — a factor commonly overlooked in parallel distributed SGD methodologies. In particular, the mixing process between any two agents is initiated solely when there is a positive policy advantage. Adhering to Theorem 2, the mixture metric is set marginally below its computed upper limit, ensuring both policy improvement and convergence.
V-B Communication Cost Analysis
Since the pulling request does not contain any actual data, its cost in the analysis can be ignored, and we only consider the policy parameters transmitted among agents to analyze the communication efficiency of RSM-MASAC.
Regarding the communication overhead for each segment request, RSM-MASAC incurs a data transmission cost of $\upsilon_{p}=\upsilon/P$ through V2V communications. Notably, we consider V2V communication as an implementation method without restricting it to specific protocols. Following the same settings as in [70], we provide illustrative examples to demonstrate the overall reduction in communication overhead. Specifically, the Collective Perception Message (CPM) defined in Collective Perception Service (CPS) of ETSI TR 103 562 [72] can encapsulate DNN parameters into the Perceived Object Containers (POCs) and propagate them among cooperating CAVs with each message carrying a payload of $4,480$ bytes, serving as segment responses in our algorithm. And the DNN parameters $\theta\in\mathbb{R}^{d}$ ’s transmission size $\upsilon$ can be regarded as $32d$ bits, commonly assumed in [62, 70, 24]. Consequently, the total communication overhead for reconstructing $\kappa$ referential policies per agent in each round $c(\upsilon,f)=\frac{N\times\upsilon\times\kappa}{8\times 1024^{3}}$ (GB), as well as $\frac{N\times\upsilon\times\kappa}{8\times 4480}$ message numbers, which is $(N-1)/\kappa$ times less than that in a fully connected communication setup [44]. Additionally, communication occurs at a periodic interval of $U$, allowing for further reduction in communication overhead by decreasing communication frequency. Moreover, by simultaneously requesting $P$ agents in parallel, RSM-MASAC benefits from the sufficient use of the bandwidth and enhances the capability to overcome possible channel degradation.
V-C Complexity Analysis
As the calculation of policy advantage $\mathbb{A}^{+}_{\pi}(\tilde{\pi})$ depends on the sample size $M$ and parameter size $|d|$, with complexity $\mathcal{O}(M|d|)$. FIM calculation with sampling approximated method using first-order gradients in (20) scales with the number of parameters, with complexity $\mathcal{O}(M|d|^{2})$. Hence, with $\kappa$ replicas, the computation complexity of CommMix is $\mathcal{O}(\kappa M|d|^{2})$ and memory complexity is $\mathcal{O}(M|d|^{2})$. However, the actual computation of the FIM is not restricted to our sampling-based method alone. Considering models with a large number of parameters, there is extensive research on FIM approximation methods, such as Diagonal Approximation [73, 74] with $\mathcal{O}(\kappa M|d|r_{\text{rank}})$, Low-Rank Approximation [75] with $\mathcal{O}(\kappa M|d|)$, etc. Thus, the complexity can be further reduced.
VI Experimental Results and Discussions
In this section, we validate the effectiveness of our proposed algorithm for CAVs speed control in IoV, highlighting its superiority compared to other methods.
VI-A Experimental Settings
We implement two simulation scenarios on Flow [76, 77], which is a traffic control benchmarking framework for mixed autonomy traffic. As illustrated in Fig. 3, the common urban traffic intersection scenario “Figure 8” and highway scenario “Merge” are selected, with the main system settings described in Table III.
(a) “Figure 8”
(b) “Merge”
Figure 3: The two scenarios for simulations on Flow. The red vehicles are DRL-driven CAVs, while the blue vehicles are the HDVs observed by the DRL-driven CAVs, and the white vehicles are the HDVs that are not observed in the state space.
TABLE III: Setting of System Parameters in two scenarios.
Parameters Definition Figure Eight Merge
Number of DRL agents $N$ $9$ $13$
Total time-steps per epoch $T$ $1,500$ $750$
Number of epochs $300$ $300$
Range of acceleration ($m/s^{2}$) $[-3,3]$ $[-1.5,1.5]$
Desired velocity per vehicle ($m/s$) $20$ $20$
Speed limit per vehicle ($m/s$) $30$ $30$
Length per time-step (s) $0.1$ $0.1$
Maximum of vehicles per hour - $2,300$
  • “Figure 8”: $14$ vehicles navigate a one-way lane shapes like a figure “8”, including $5$ emulated HDVs controlled by Simulation of Urban MObility (SUMO) with Intelligent Driver Model (IDM) [78], and $9$ IRL-controlled CAVs maintaining dedicated links to update their parameters through the V2V channel. At the lane’s intersection, each CAV adjusts its acceleration to traverse efficiently, aiming to boost the traffic flow’s average speed.
  • “Merge”: A highway on-ramp merging scenario with vehicle flow at $2,300$ per hour, including a maximum of $2,200$ vehicles on the main road and $100$ vehicles on the ramp. Within each epoch, $13$ vehicles are randomly chosen to instantiate the DRL-based controllers, aiming to manage collision avoidance and congestion at merge points. The simulation settings closely align with those of the first scenario.
Both two scenarios are modified to assign the limited partial observation of the global environment as the state of each CAV, including the position and speed of its own, the vehicle ahead and behind. Only CAVs can execute the V2V end-to-end communication, and the connectivity of V2V links at time $t$ depends on the CAVs’ position and communication range, which are both extracted using the TraCI simulator in our experiments. A communication range of $90$m is adopted in interactions “Figure 8” and $400$m in highway “Merge”. This is according to the conclusion in [79] that communication range at intersections will be extremely reduced compared with the conventional scenarios [80], and relevant service requirements[81]. Unless otherwise stated, our experiments are based on the common MARL V2V lossless ideal communication premise [30, 31, 32, 33]. Meanwhile, each CAV’s action is a continuous variable representing the speed acceleration or deceleration and is normalized between $[-1,1]$.
TABLE IV: Hyper-parameters.
Hyper-parameters Symbol Value
Replay buffer size $|\mathcal{D}|$ $10^{5}$
Batch size $\Phi$ $256$
Number of samples to evaluate $\mathbb{A}^{+}_{\pi}(\tilde{\pi})$ and $F(\theta)$ $M$ $50$
Learning rate of actor network $\eta_{\pi}$ $10^{-4}$
Learning rate of critic network $\eta_{Q}$ $3\times 10^{-3}$
Learning rate of temperature parameter $\eta_{\alpha}$ $3\times 10^{-4}$
Discount factor $\gamma$ $0.99$
Target smoothing coefficient $\varrho$ $10^{-3}$
Delayed policy update intervals $e$ $10$
Communication intervals $U$ $8$
Number of segments $P$ 4
Number of model replicas $\kappa$ 3
(a) “Figure 8”
(b) “Merge”
Figure 4: Performance of IRL without communication.
(a) “Figure 8”
(b) “Merge”
Figure 5: Performance comparison of different methods.
In order to reduce the occurrence of collisions and promote the traffic flow to the maximum desired speed, we take the normalized average speed of all vehicles at each timestep as the individual reward in each scenario [^2] Notably, we assume complete knowledge of individual vehicle speeds at each vehicle here. Beyond the scope of this paper, some value-decomposition methods like [82] can be further leveraged to derive a decomposed reward, so as to loosen such a strict requirement., which is assigned to each training vehicle after its action is performed. In addition, the current epoch will be terminated once a collision occurs or the max length of step $T$ in an epoch is reached.
We perform tests every $10$ epochs and calculate the average reward based on the accumulated rewards from $5$ independent testing epochs. Besides, all results are produced by taking the average of $5$ independent simulations. Moreover, as for the baseline, we take the vehicles controlled by the Flow IDM [78], which belongs to a typical car-following model incorporating extensive prior knowledge and indicates the pinnacle of performance achievable by the best centralized federated MARL algorithms [27]. Furthermore, the principal hyper-parameters used in simulations are listed in Table IV.
VI-B Evaluation Metrics
Apart from the average reward, we adopt some additional metrics to extensively evaluate the communication efficiency of RSM-MASAC.
  • We denote the total count of reconstructed referential policies (i.e., all model replicas) as $\rho_{\text{total}}$. The number of effectively reconstructed referential policies, those contributing to the mixing process, is represented by $\rho_{\text{ef}}$. The mixing rate $\rho_{r}=\rho_{\text{ef}}/\rho_{\text{total}}$ thus reflects the usage rate of reconstructed policies.
  • We use $\psi$ to indicate the overall communication overhead (in terms of $\upsilon$) in an epoch, and $C_{0}$ denotes the number of communication rounds. Therefore, the communication overheads equal $\psi=C_{0}\times c(\upsilon,f)=\frac{\rho_{\text{total}}\times\upsilon}{8\times 1024^{3}}$ (GB).
Figure 6: The effect of segmentation on mixing rate.
(a) Communication times.
(b) Effective mixture times.
Figure 7: Heatmap comparison of communication times and effective mixture times under “Figure 8”.
VI-C Simulation Results
TABLE V: Results of Average Reward & Communication Efficiency under “Merge” scenario.
Method $U$ $P$ $\kappa$ Average Reward $\rho_{\text{total}}$ $\psi$ (GB) $\rho_{\text{ef}}$ $\rho_{r}$
Avg- MAPPO $8$ $4$ $3$ $0.5940\pm 0.0204$ $24,381$ $6.185$ $24,381$ $100\%$
MASAC $0.6584\pm 0.0107$ $43,797$ $11.111$ $43,797$ $100\%$
RSM- MAPPO $0.6173\pm 0.0097$ $24,381$ $6.185$ $10,294$ $42.22\%$
MASAC $1$ $0.6648\pm 0.0065$ $14,599$ $3.704$ $7,938$ $54.37\%$
$5$ $0.6745\pm 0.0120$ $72,995$ $18.518$ $33,484$ $45.87\%$
$7$ $0.6721\pm 0.0072$ $102,102$ $25.902$ $47,122$ $46.15\%$
$3$ $0.6658\pm 0.0063$ $43,797$ $11.111$ $22,661$ $51.74\%$
$1$ $0.6654\pm 0.0048$ $43,797$ $11.111$ $21,024$ $48.00\%$
$2$ $0.6677\pm 0.0064$ $43,797$ $11.111$ $21,083$ $48.14\%$
$6$ $0.6656\pm 0.0059$ $43,797$ $11.111$ $22,128$ $50.52\%$
$72$ $4$ $0.6464\pm 0.0292$ $4,875$ $1.237$ $2,345$ $48.10\%$
$144$ $0.6455\pm 0.0486$ $2,457$ $0.623$ $1,046$ $42.57\%$
Beforehand, we present the performance of Independent SAC (ISAC) and Independent PPO (IPPO) for MARL without any information sharing in Fig. 4, so as to highlight the critical role of inter-agent communication in decentralized cooperative MARL and facilitate subsequent discussions. Fig. 4 reveals that the learning process of these non-cooperative algorithms achieves unstable average reward with greater variance, and suffers from convergence issues even until $300$ epochs. Whereas, when a communication-assisted mixing phase involving the exchange of policy parameters is incorporated, as shown in Fig. 5, the learning process within the multi-agent environment is remarkably enhanced in terms of stability and efficiency. In other words, consistent with our previous argument, integrating DFL into IRL significantly improves training efficiency and ensures learning stability.
In Fig. 5, we also compare the performance of Regulated Segment Mixture (RSM) and that of the direct, unselective Averaging (Avg) method, as mentioned in Section III. It can be observed in Fig. 5, in terms of convergence speed and stability, the simple average mixture method is somewhat inferior. Particularly in more complex scenarios, as in “Merge”, which involves a greater number of RL agents and denser traffic flow compared to the “Figure 8”, the disparity between these two methods becomes more pronounced. The larger variance in the average reward under the average mixture method suggests a more unstable parameter mixing process. Therefore, it validates the effectiveness of RSM and supports the derived theoretical results for selecting useful reference policies and assigning appropriate mixing weights.
On the other hand, Fig. 5 offers a further evaluation of the RSM’s performance on both MASAC and MAPPO. In particular, proposed in our previous work [1], RSM-MAPPO utilizes PPO under traditional MARL framework as the independent local learning algorithm, making it a special case of RSM-MASAC with $\alpha=0$ and a different policy advantage estimation under our more general reanalyses. Evidently, in both scenarios, the SAC-based algorithm demonstrates faster convergence and overall superior performance compared to the PPO-based algorithm under the same learning rate. We believe this is primarily attributed to the differences in exploration mechanism and sample efficiency. Specifically, benefiting from the introduced entropy item under MERL framework, SAC encourages policies to better explore environments, thus capably avoiding local optima and potentially learning faster.
(a) the number of segments $P$
(b) the number of replicas $\kappa$
Figure 8: The performance of RSM-MASAC with respect to different hyperparameters ((a) the number of segments $P$ and (b) the number of replicas $\kappa$), based on the testing episodes after $150$, $160$, $\cdots$, $250$ training epochs under “Merge”.
Figure 9: Impact of different target smoothing coefficient $\varrho$ under “Merge”.
Figure 10: Double vs single $Q$ network under “Merge”.
Next, we evaluate the impact of regulated segmentation. Fig. 7 shows that the incorporation of segmentation (i.e., $P=4$) also leads to an improvement in the successful mixing rate $\rho_{\text{r}}$ of the policy, apart from better utilizing the available bandwidth as discussed in Section V-B. The improvement lies in the introduction of a certain level of randomness, consistent with the idea to encourage exploration by adding an entropy term in MERL. But this randomness is also regulated by the policy parameter mixing theorem in Theorem 2 without compromising performance. To embody the selection process of reconstructed referential policies, we further present the heatmap comparison between communication times and successful mixture times in Fig. 7. It can be observed from Fig. 7(a) that vehicles communicate frequently with their neighboring vehicles, but since not all communication packets (i.e., model segments) are utilized for improving local policy performance, it leads to the uneven and asymmetrical heatmap for the successful mixture of segments in Fig. 7(b). Our derived performance improvement bound and practical parameter mixture metric regulate these segments, demonstrating that the data from different agents contribute variably to the mixture.
In Fig. 8, we further discuss the effects of the number of segments $P$ and replicas $\kappa$ under RSM-MASAC. Considering that RSM-MASAC generally converges from $250$ to $300$ testing epochs, the results in the boxplots of Fig. 8 are derived from testing epochs ranging from $150$ to $250$, during which RSM-MASAC experiences more fluctuation and slower convergence, so as to better illustrate the performance during training. Specifically, Fig. 8(a) displays the distribution of average reward, where each box represents the InterQuartile Range (IQR). The lower and upper edges of each box denote the first and third quartiles, respectively. Besides, the line inside the box signifies the median average reward for each segment number. Observing Fig. 8(a), a trend emerges showing that the median average reward subtly increases with the growing number of segments. Besides, Fig. 8(b) shows the median average reward slightly increases along with larger $\kappa$ values, suggesting a trend that more replicas may produce with a higher reward or a higher probability of exploring a better policy, at some cost of a heavier perturbation of the learning process.
Figure 11: Impact of different communication intervals $U$ under “Merge”.
Figure 12: Performance improvement of continually fine-tuning under “Figure 8”.
Figure 13: Performance comparison of different PRR in communication under “Figure 8”.
In addition, we take the average reward within $250-300$ testing epochs as the final converged performance, and summarize corresponding experimental results and details about communication overhead in Table V, which takes the V2V communication examples’ settings in V-B. We can find that with $P$ increasing, the mixing rate $\rho_{r}$ also increases first, but slightly decreases when $P=6$, since the aggregation target of reconstructed policy parameters for an over-large $P$ might be mottled and lose the integrity. Besides, a larger $\kappa$ gives the increase of the communication overhead $\psi$, while the final performance gains, especially for $\kappa=5$, $\kappa=7$, are limited, due to that the algorithm has converged to a relatively good policy as in centralized FMARL [27].
In addition, we conduct more ablation studies to evaluate the design of $Q$ network on the learning process. In Fig. 10, we analyze the impact of different target smoothing coefficients $\varrho$ on the learning process. It can be observed that a larger value of the target smoothing coefficient, such as $10^{-2}$, can significantly accelerate the learning, but it also introduces disturbances, leading to non-steady learning. On the other hand, a smaller value like $10^{-4}$ or $10^{-5}$ noticeably slows down the learning speed, affecting the performance of the algorithm. Hence, we take $\varrho=10^{-3}$ as the default value. Moreover, in Fig. 10, we also evaluate the adoption of dual $Q$ networks. It can be observed that employing dual $Q$ networks indeed accelerates the learning process.
Furthermore, we also discuss the impact of communication intervals on the learning process, as shown in Fig. 11. It is evident that larger communication intervals result in reduced learning speeds and significant instability in the learning process.
Meanwhile, we also investigate the effect of model fine-tuning in decentralized IoV speed control settings. Specifically, in an environment full of $14$ DRL-controlled CAVs, we compare the performance between a continually fine-tuning policy and a well-trained, convergent policy (i.e., RSM-MASAC in Fig. 5) to execute inference only (i.e., test only in DRL description). The results in Fig. 13 show that the fine-tuned policy in new conditions can reduce the occurrence of collisions at the beginning and continuously achieve better performance than just testing. Furthermore, to investigate the impact of possible packet loss when communicating on V2V links, we provide the result for different Packet Reception Rate (PRR) in Fig. 13. It can be seen that a packet loss of less than $20\%$ has a trivial effect on performance, demonstrating that DNN parameter transmission in VFL can yield good robustness in most real communication environments.
VII Conclusions
In this paper, we have proposed a communication-efficient algorithm RSM-MASAC as a promising solution to enhance communication efficiency and policy collaboration in distributed MARL, particularly in the context of dynamic IoV environments. By delving into the policy parameter mixture function, RSM-MASAC has provided a novel means to leverage and boost the effectiveness of distributed multi-agent collaboration. In particular, RSM-MASAC has successfully transformed the classical means of complete parameter exchange into segment-based request and response, which significantly facilitates the construction of multiple referential polices and simultaneously captures enhanced learning diversity. Moreover, in order to avoid performance-harmful parameter mixture, RSM-MASAC has leveraged a theory-established regulated mixture metric, and selects the contributive referential policies with positive relative policy advantage only. Finally, extensive simulations in the mixed-autonomy traffic control scenarios have demonstrated the effectiveness of the proposed approach.
Appendix: Proofs
VII-A Proof of Soft Policy Improvement
Lemma 1.
(Soft policy improvement) Let $\pi_{\text{old}}$ and $\pi_{\text{new}}$ be the optimizer of the minimization problem defined in (5). Then $Q^{\pi_{\text{new}}}(s,a)\geq Q^{\pi_{\text{old}}}(s,a),\forall s,a$ with $|\mathcal{A}|<\infty$.
Proof.
This proof is a direct application of soft policy improvement [42, 43]. We leave the proof here for completeness.
Let $\pi_{\text{old}}\in\Pi$ and $Q^{\pi},V^{\pi}$ is the corresponding soft state-action value and soft state value, respectively. And $\pi_{\text{new}}$ is defined as:
$\pi_{\mathrm{new}}$ $=\arg\min_{\pi\in\prod}\mathrm{D}_{\mathrm{KL}}\left(\pi(\cdot|s)\|\frac{\exp\left(\frac{1}{\alpha}Q^{\pi_{\text{old}}}({s},\cdot)\right)}{Z^{\pi_{\text{old}}}(s)}\right)$
$=\arg\min_{\pi\in\prod}J_{\pi_{\text{old}}}(\pi(\cdot|s))$
Since we can always choose $\pi_{\text{new}}=\pi_{\text{old}}\in\Pi$, there must be $J_{\pi_{\text{old}}}(\pi_{\text{new}}(\cdot|s))\leq J_{\pi_{\text{old}}}(\pi_{\text{old}}(\cdot|s))$. Hence
$\mathbb{E}_{{a}\sim{\pi_{\mathrm{new}}}}\left[\frac{1}{\alpha}Q^{\pi_{\text{old}}}({s},{a})-\log\pi_{\text{new}}(a|s)-\log Z^{\pi_{\text{old}}}(s)\right]$
$\geq\mathbb{E}_{{a}\sim\pi_{\text{old}}}\left[\frac{1}{\alpha}Q^{\pi_{\text{old}}}({s},{a})-\log\pi_{\text{old}}(a|s)-\log Z^{\pi_{\text{old}}}(s)\right]$
as partition function $Z$ depends only on the state, the inequality reduces to a form of the sum of entropy and value with one-step look-ahead
$\mathbb{E}_{{a}\sim{\pi_{\mathrm{new}}}}$ $\left[Q^{\pi_{\text{old}}}({s},{a})\right]+\alpha H({\pi_{\mathrm{new}}}(\cdot|{s}))$
$\geq\mathbb{E}_{{a}\sim\pi_{\text{old}}}\left[Q^{\pi_{\text{old}}}({s},{a})\right]+\alpha H(\pi_{\text{old}}(\cdot|{s}))=V_{\pi_{\text{old}}}(s)$
And according to the definition of the soft $Q$-value in section II, we can get that
$Q^{\pi_{\text{old}}}(s,a)$
$=$ $\mathop{\mathbb{E}}_{\begin{subarray}{c}s_{1}\end{subarray}}\left[r_{0}\!+\!\gamma\!\left(\!\alpha H(\pi_{\text{old}}(\cdot|s_{1}))\!+\!\!\mathop{\mathbb{E}}_{\begin{subarray}{c}a_{1}\sim\pi_{\text{old}}\end{subarray}}[Q^{\pi_{\text{old}}}(s_{1},a_{1})]\!\right)\!\right]$ (21)
$\leq$ $\mathop{\mathbb{E}}_{\begin{subarray}{c}s_{1}\end{subarray}}\left[r_{0}\!+\!\gamma\!\left(\!\alpha H({\pi_{\text{new}}}(\cdot|s_{1}))\!+\!\mathop{\mathbb{E}}_{\begin{subarray}{c}a_{1}\sim{\pi_{\text{new}}}\end{subarray}}[Q^{\pi_{\text{old}}}(s_{1},a_{1})]\!\right)\!\right]$
$=$ $\mathop{\mathbb{E}}_{\begin{subarray}{c}s_{1}\\ a_{1}\sim{\pi_{\text{new}}}\end{subarray}}\Big{[}r_{0}+\gamma\left(\alpha H({\pi_{\text{new}}}(\cdot|s_{1}))+r_{1}\right)\Big{]}$
$\quad+\gamma^{2}\mathop{\mathbb{E}}_{\begin{subarray}{c}s_{2}\end{subarray}}\left[\alpha H(\pi_{\text{old}}(\cdot|s_{2}))+\mathop{\mathbb{E}}_{\begin{subarray}{c}a_{2}\sim\pi_{\text{old}}\end{subarray}}[Q^{\pi_{\text{old}}}(s_{2},a_{2})]\right]$
$\leq$ $\mathop{\mathbb{E}}_{\begin{subarray}{c}s_{1}\\ a_{1}\sim{\pi_{\text{new}}}\end{subarray}}\left[r_{0}+\gamma\left(\alpha H({\pi_{\text{new}}}(\cdot|s_{1}))+r_{1}\right)\right]$
$\quad+\gamma^{2}\mathop{\mathbb{E}}_{\begin{subarray}{c}s_{2}\end{subarray}}\left[\alpha H({\pi_{\text{new}}}(\cdot|s_{2}))+\mathop{\mathbb{E}}_{\begin{subarray}{c}a_{2}\sim{\pi_{\text{new}}}\end{subarray}}[Q^{\pi_{\text{old}}}(s_{2},a_{2})]\right]$
$=$ $\!\mathop{\mathbb{E}}_{\begin{subarray}{c}s_{1}\\ a_{1}\sim{\pi_{\text{new}}}\\ s_{2}\end{subarray}}\!\!\left[r_{0}\!+\!\gamma\!\left(\alpha H({\pi_{\text{new}}}(\cdot|s_{1}))\!+\!r_{1}\right)\!+\!\gamma^{2}\!\left(\alpha H({\pi_{\text{new}}}(\cdot|s_{2}))\!+\!r_{2}\right)\right]$
$\quad+\gamma^{3}\mathop{\mathbb{E}}_{\begin{subarray}{c}s_{3}\end{subarray}}\left[\alpha H({\pi_{\text{new}}}(\cdot|s_{3}))+\mathop{\mathbb{E}}_{\begin{subarray}{c}a_{3}\sim{\pi_{\text{new}}}\end{subarray}}[Q^{\pi_{\text{old}}}(s_{3},a_{3})]\right]$
[…]
$\leq$ $\mathop{\mathbb{E}}\limits_{\tau_{1}\sim\pi_{\text{new}}}\left[r_{0}+\sum_{t=1}^{\infty}\gamma^{t}\big{(}H({\pi_{\text{new}}}(\cdot|s_{t}))+r_{t}\big{)}\right]$
$=$ $Q^{\pi_{\text{new}}}(s,a)$
VII-B Proof of Theorem 1
Before proving the theorem, we first introduce several important lemmas, on which the proofs of the proposed theorems are built.
Lemma 2.
$\mathbb{E}_{a\sim\pi}\left[A_{\pi}(s,a)\right]=-\alpha H(\pi(\cdot|s))$
Proof.
$\mathbb{E}_{a\sim\pi}\left[A_{\pi}(s,a)\right]$
$=$ $\sum_{a}\pi(a|s)A_{\pi}(s,a)$
$=$ $\sum_{a}\pi(a|s)\left[Q_{\pi}(s,a)-V_{\pi}(s)\right]$
$=$ $\sum_{a}\pi(a|s)Q_{\pi}(s,a)-V_{\pi}(s)$
=$\stackrel{{\scriptstyle\mathrm{(a)}}}{{=}}$ $\sum_{a}\pi(a|s)Q_{\pi}(s,a)-\sum_{a}\pi(a|s)\left[Q_{\pi}(s,a)-\alpha\log\pi(a|s)\right]$
$=$ $-\alpha H(\pi(\cdot|s))$
where the equality $(a)$ is according to (4). ∎
Lemma 3.
$\mathop{\mathbb{E}}\limits_{a\sim\pi_{\text{mix}}}\left[A_{\pi}(s,a)\right]=\beta\mathop{\mathbb{E}}\limits_{a\sim\tilde{\pi}}\left[A_{\pi}(s,a)\right]-(1-\beta)\alpha H(\pi(\cdot|s))$
Proof.
$\mathbb{E}_{a\sim\pi_{\text{mix}}}\left[A_{\pi}(s,a)\right]$
$=$ $\sum_{a}\pi_{\text{mix}}(a|s)A_{\pi}(s,a)$
=$\stackrel{{\scriptstyle\mathrm{(a)}}}{{=}}$ $\sum_{a}\left[(1-\beta)\pi(a|s)+\beta\tilde{\pi}(a|s)\right]A_{\pi}(s,a)$
$=$ $\beta\sum_{a}\tilde{\pi}(a|s)A_{\pi}(s,a)+(1-\beta)\sum_{a}{\pi}(a|s)A_{\pi}(s,a)$
=$\stackrel{{\scriptstyle\mathrm{(b)}}}{{=}}$ $\beta\mathop{\mathbb{E}}\limits_{a\sim\tilde{\pi}}\left[A_{\pi}(s,a)\right]-(1-\beta)\alpha H(\pi(\cdot|s))$
where the equalities $(a)$ and $(b)$ are due to (12) and Lemma 2, respectively. ∎
Lemma 4.
$\eta(\pi_{\text{mix}})-\eta(\pi)$
$=$ $\sum_{t=0}^{\infty}\gamma^{t}\left[\mathbb{E}_{\begin{subarray}{c}s\sim P(s_{t};\pi_{\text{mix}})\\ a\sim\pi_{\text{mix}}\end{subarray}}[A_{\pi}(s,a)+\alpha H(\pi_{\text{mix}}(\cdot|s))]\right]$
Proof.
First note that according to the definition of advantage function and soft bellman equation in (2) and (4), we can get $A_{\pi}(s_{t},a_{t})=\mathbb{E}_{s_{t+1}\sim P(\cdot|s_{t},a_{t})}[r_{t}+\gamma V_{\pi}(s_{t+1})-V_{\pi}(s_{t})]$. Therefore,
$\mathop{\mathbb{E}}\limits_{\tau_{0}\sim\pi_{\text{mix}}}\left[\sum_{t=0}^{\infty}\gamma^{t}\left[A_{\pi}(s_{t},a_{t})\right]\right]$
$=$ $\mathop{\mathbb{E}}\limits_{\tau_{0}\sim\pi_{\text{mix}}}\left[\sum_{t=0}^{\infty}\gamma^{t}\left[r_{t}+\gamma V_{\pi}(s_{t+1})-V_{\pi}(s_{t})\right]\right]$
$=$ $\mathop{\mathbb{E}}\limits_{\tau_{0}\sim\pi_{\text{mix}}}\Bigg{[}\sum_{t=0}^{\infty}\gamma^{t}r_{t}+(\gamma V_{\pi}(s_{1})-V_{\pi}(s_{0})$
$\quad\quad+\gamma^{2}V_{\pi}(s_{2})-\gamma V_{\pi}(s_{1})+\cdots)\Bigg{]}$
$=$ $\mathbb{E}_{\tau_{0}\sim\pi_{\text{mix}}}\left[\sum_{t=0}^{\infty}\gamma^{t}r_{t}\right]-\mathbb{E}_{s_{0}}\left[V_{\pi}(s_{0})\right]$
[…]
$\quad\left.-\sum_{t=0}\gamma^{t}[\alpha H(\pi_{\text{mix}}(\cdot|s_{t}))]\right]-\eta(\pi)$
$=$ $\eta(\pi_{\text{mix}})-\eta(\pi)-\mathbb{E}_{\tau_{0}\sim\pi_{\text{mix}}}\left[\sum_{t=0}^{\infty}\gamma^{t}[\alpha H(\pi_{\text{mix}}(\cdot|s_{t}))]\right]$
Also we have
$\mathop{\mathbb{E}}\limits_{\tau_{0}\sim\pi_{\text{mix}}}\left[\sum_{t=0}^{\infty}\gamma^{t}\left[A_{\pi}(s_{t},a_{t})+\alpha H(\pi_{\text{mix}}(\cdot|s_{t}))\right]\right]$
$=$ $\sum_{t=0}^{\infty}\gamma^{t}\left[\mathbb{E}_{\begin{subarray}{c}s\sim P(s_{t};\pi_{\text{mix}})\\ a\sim\pi_{\text{mix}}\end{subarray}}[A_{\pi}(s,a)+\alpha H(\pi_{\text{mix}}(\cdot|s))]\right]$
Hence we can derive the formula for Lemma 4. ∎
Lemma 5.
For any given state $s$, we can decompose the entropy of the mixed policy as
$H(\pi_{\text{mix}}(\cdot|s))$ $=\mathrm{D}_{\mathrm{JS}}^{\beta}(\tilde{\pi}(\cdot|s)\|\pi(\cdot|s))+\beta H(\tilde{\pi}(\cdot|s))$
$\qquad+(1-\beta)H(\pi(\cdot|s))$
Proof.
By definition,
$H(\pi_{\text{mix}}(\cdot|s))$
$=$ $-\sum_{a}\Big{\{}[(1-\beta)\pi(a|s)+\beta\tilde{\pi}(a|s)]$
$\qquad\qquad\cdot\log[(1-\beta)\pi(a|s)+\beta\tilde{\pi}(a|s)]\Big{\}}$
$=$ $\sum_{a}\beta\left[\tilde{\pi}(a|s)\log\frac{\tilde{\pi}(a|s)}{(1-\beta)\pi(a|s)+\beta\tilde{\pi}(a|s)}\right]$
$\quad+\sum_{a}(1-\beta)\left[{\pi}(a|s)\log\frac{{\pi}(a|s)}{(1-\beta)\pi(a|s)+\beta\tilde{\pi}(a|s)}\right]$
$\quad-\sum_{a}\left[\beta\tilde{\pi}(a|s)\log\tilde{\pi}(a|s)\right]$
$\quad-\sum_{a}\left[(1-\beta){\pi}(a|s)\log{\pi}(a|s)\right]$
$=$ $\mathrm{D}_{\mathrm{JS}}^{\beta}(\tilde{\pi}(\cdot|s)\|\pi(\cdot|s))+\beta H(\tilde{\pi}(\cdot|s))+(1-\beta)H(\pi(\cdot|s))$
Lemma 6.
$\mathop{\mathbb{E}}_{\begin{subarray}{c}{s\sim P(s_{t};\pi_{\text{mix}})}\\ {a\sim\pi_{\text{mix}}}\end{subarray}}\left[A_{\pi}(s,a)+\alpha H(\pi_{\text{mix}}(\cdot|s))\right]$
$\geq$ $\quad\beta\mathop{\mathbb{E}}_{\begin{subarray}{c}s\sim P(s_{t};\pi)\\ a\sim\tilde{\pi}\end{subarray}}\left[A_{\pi}(s,a)+\alpha H(\tilde{\pi}(\cdot|s))\right]-2\beta\rho_{t}\varepsilon$
$\quad+\alpha\mathop{\mathbb{E}}\limits_{s\sim P(s_{t};\pi_{\text{mix}})}\left[D_{\text{JS}}^{\beta}(\tilde{\pi}(\cdot|s)\|\pi(\cdot|s))\right]$
where $\varepsilon=\max_{s}|\mathbb{E}_{a\sim\tilde{\pi}}[A_{\pi}(s,a)+\alpha H(\tilde{\pi}(\cdot|s))]|$, and $\rho_{t}=1-(1-\beta)^{t}$.
Proof.
$\mathop{\mathbb{E}}_{\begin{subarray}{c}{s\sim P(s_{t};\pi_{\text{mix}})}\\ {a\sim\pi_{\text{mix}}}\end{subarray}}\left[A_{\pi}(s,a)+\alpha H(\pi_{\text{mix}}(\cdot|s))\right]$
=$\stackrel{{\scriptstyle\mathrm{(a)}}}{{=}}\mathop{\mathbb{E}}\limits_{s\sim P(s_{t};\pi_{\text{mix}})}\Bigg{[}\beta\sum_{a}\tilde{\pi}(a|s)A_{\pi}(s,a)-(1-\beta)\alpha H(\pi(\cdot|s))$
$\quad+\alpha\mathrm{D}_{\mathrm{JS}}^{\beta}(\tilde{\pi}(\cdot|s)\|\pi(\cdot|s))\!+\!\beta\alpha H(\tilde{\pi}(\cdot|s))\!+\!(1\!-\!\beta)\alpha H(\pi(\cdot|s))\Bigg{]}$
$=\beta\mathop{\mathbb{E}}\limits_{s\sim P(s_{t};\pi_{\text{mix}})}\left[\sum_{a}\tilde{\pi}(a|s)\Big{[}A_{\pi}(s,a)+\alpha H(\tilde{\pi}(\cdot|s))\Big{]}\right]$
$\quad+\alpha\mathop{\mathbb{E}}\limits_{s\sim P(s_{t};\pi_{\text{mix}})}\left[\mathrm{D}_{\mathrm{JS}}^{\beta}(\tilde{\pi}(\cdot|s)\|\pi(\cdot|s))\right]$
where the equality $(a)$ is according to Lemma 3 and Lemma 5, and the mixed policy is taken as a mixture of the policy $\pi$ and the referential policy $\tilde{\pi}$ received from others. In other words, to sample from $\pi_{\text{mix}}$, we first draw a Bernoulli random variable, which tells us to choose $\pi$ with probability $(1-\beta)$ and choose $\tilde{\pi}$ with probability $\beta$. Let $c_{t}$ be the random variable that indicates the number of times $\tilde{\pi}$ was chosen before time $t$. $P(s_{t};\pi)$ is the distribution over states at time $t$ while following $\pi$. We can condition on the value of $c_{t}$ to break the probability distribution into two pieces, with $P(c_{t}=0)=(1-\beta)^{t}$, and $\rho_{t}=P(c_{t}\geq 1)=1-(1-\beta)^{t}$. Thus, we can get (22) on Page 22.
$\beta\mathop{\mathbb{E}}\limits_{s\sim P(s_{t};\pi_{\text{mix}})}\left[\sum_{a}\tilde{\pi}(a|s)\left[A_{\pi}(s,a)+\alpha H(\tilde{\pi}(\cdot|s))\right]\right]$
$=$ $\beta(1-\rho_{t})\!\mathop{\mathbb{E}}\limits_{s\sim P(s_{t}|c_{t}=0;\pi_{\text{mix}})}\!\left[\sum_{a}\tilde{\pi}(a|s)\left[A_{\pi}(s,a)+\alpha H(\tilde{\pi}(\cdot|s))\right]\right]+\beta\rho_{t}\mathop{\mathbb{E}}\limits_{s\sim P(s_{t}|c_{t}\geq 1;\pi_{\text{mix}})}\left[\sum_{a}\tilde{\pi}(a|s)\left[A_{\pi}(s,a)+\alpha H(\tilde{\pi}(\cdot|s))\right]\right]$
$=$ $\beta\!\mathop{\mathbb{E}}\limits_{s\sim P(s_{t}|c_{t}=0;\pi_{\text{mix}})}\!\left[\sum_{a}\tilde{\pi}(a|s)\left[A_{\pi}(s,a)+\alpha H(\tilde{\pi}(\cdot|s))\right]\right]-\beta\rho_{t}\mathop{\mathbb{E}}\limits_{s\sim P(s_{t}|c_{t}=0;\pi_{\text{mix}})}\left[\sum_{a}\tilde{\pi}(a|s)\left[A_{\pi}(s,a)+\alpha H(\tilde{\pi}(\cdot|s))\right]\right]$ (22)
$\quad+\beta\rho_{t}\mathop{\mathbb{E}}\limits_{s\sim P(s_{t}|c_{t}\geq 1;\pi_{\text{mix}})}\left[\sum_{a}\tilde{\pi}(a|s)\left[A_{\pi}(s,a)+\alpha H(\tilde{\pi}(\cdot|s))\right]\right]$
Recalling the definition $\varepsilon=\max_{s}\big{|}\mathbb{E}_{a\sim\tilde{\pi}}[A_{\pi}(s,a)+\alpha H(\tilde{\pi}(\cdot|s))]\big{|}$, we have
$\mathop{\mathbb{E}}\limits_{s\sim P(s_{t}|c_{t}=0;\pi_{\text{mix}})}\left[\sum_{a}\tilde{\pi}(a|s)\left[A_{\pi}(s,a)+\alpha H(\tilde{\pi}(\cdot|s))\right]\right]\leq\varepsilon$
$\mathop{\mathbb{E}}\limits_{s\sim P(s_{t}|c_{t}\geq 1;\pi_{\text{mix}})}\left[\sum_{a}\tilde{\pi}(a|s)\left[A_{\pi}(s,a)+\alpha H(\tilde{\pi}(\cdot|s))\right]\right]\geq-\varepsilon$
Therefore, (22) can be re-organized as
$\beta\mathop{\mathbb{E}}\limits_{s\sim P(s_{t};\pi_{\text{mix}})}\left[\sum_{a}\tilde{\pi}(a|s)\left[A_{\pi}(s,a)+\alpha H(\tilde{\pi}(\cdot|s))\right]\right]$
$\geq$ $\beta\!\mathop{\mathbb{E}}\limits_{s\sim P(s_{t}|c_{t}=0;\pi_{\text{mix}})}\!\left[\sum_{a}\tilde{\pi}(a|s)\left[A_{\pi}(s,a)+\alpha H(\tilde{\pi}(\cdot|s))\right]\right]$
$\quad-2\beta\rho_{t}\varepsilon$
$=$ $\beta\!\mathop{\mathbb{E}}\limits_{s\sim P(s_{t};\pi)}\!\left[\sum_{a}\tilde{\pi}(a|s)\left[A_{\pi}(s,a)+\alpha H(\tilde{\pi}(\cdot|s))\right]\right]\!-\!2\beta\rho_{t}\varepsilon$
in which ${P(s_{t}|c_{t}=0;\pi_{\text{mix}})}={P(s_{t};\pi)}$. ∎
Next, we are ready to prove Theorem 1.
Proof.
According to Lemma 4 and Lemma 6, we have
$\eta(\pi_{\text{mix}})-\eta(\pi)$
$=$ $\sum_{t=0}^{\infty}\gamma^{t}\left[\mathbb{E}_{\begin{subarray}{c}s\sim P(s_{t};\pi_{\text{mix}})\\ a\sim\pi_{\text{mix}}\end{subarray}}[A_{\pi}(s,a)+\alpha H(\pi_{\text{mix}}(\cdot|s))]\right]$
$\geq$ $\beta\sum_{t=0}^{\infty}\gamma^{t}\!\mathop{\mathbb{E}}_{\begin{subarray}{c}s\sim P(s_{t};\pi)\\ a\sim\tilde{\pi}\end{subarray}}\left[A_{\pi}(s,a)+\alpha H(\tilde{\pi}(\cdot|s))\right]-2\beta\varepsilon\sum_{t=0}^{\infty}\gamma^{t}\rho_{t}$
$\quad+\alpha\sum_{t=0}^{\infty}\gamma^{t}\mathop{\mathbb{E}}\limits_{s\sim P(s_{t};\pi_{\text{mix}})}\left[\mathrm{D}_{\mathrm{JS}}^{\beta}(\tilde{\pi}(\cdot|s)\|\pi(\cdot|s))\right]$
$=$ $\beta\mathop{\mathbb{E}}_{\begin{subarray}{c}s\sim d_{\pi}\\ a\sim\tilde{\pi}\end{subarray}}\left[A_{\pi}(s,a)+\alpha H(\tilde{\pi}(\cdot|s))\right]-2\beta\varepsilon\sum_{t=0}\gamma^{t}[1-(1-\beta)^{t}]$
$\quad+\alpha\mathop{\mathbb{E}}_{s\sim d_{\pi_{\text{mix}}}}\left[\mathrm{D}_{\mathrm{JS}}^{\beta}(\tilde{\pi}(\cdot|s)\|\pi(\cdot|s))\right]$
$=$ $\beta\mathop{\mathbb{E}}_{\begin{subarray}{c}s\sim d_{\pi}\\ a\sim\tilde{\pi}\end{subarray}}\left[A_{\pi}(s,a)+\alpha H(\tilde{\pi}(\cdot|s))\right]-\frac{2\gamma\varepsilon\beta^{2}}{(1\!-\!\gamma)(1\!-\!\gamma(1-\beta))}$
$\quad+\alpha\mathop{\mathbb{E}}_{s\sim d_{\pi_{\text{mix}}}}\left[\mathrm{D}_{\mathrm{JS}}^{\beta}(\tilde{\pi}(\cdot|s)\|\pi(\cdot|s))\right]$
$\geq$ $\beta\mathop{\mathbb{E}}_{\begin{subarray}{c}s\sim d_{\pi}\\ a\sim\tilde{\pi}\end{subarray}}\left[A_{\pi}(s,a)+\alpha H(\tilde{\pi}(\cdot|s))\right]-\frac{2\gamma\varepsilon\beta^{2}}{(1-\gamma)^{2}}$
$\quad+\alpha\mathop{\mathbb{E}}_{s\sim d_{\pi_{\text{mix}}}}\left[\mathrm{D}_{\mathrm{JS}}^{\beta}(\tilde{\pi}(\cdot|s)\|\pi(\cdot|s))\right]$
We have the theorem. ∎
VII-C Derivation of (19)
Lemma 7.
$\mathbb{A}^{+}_{\pi}(\tilde{\pi})$ $\approx\mathbb{E}_{s_{t},a_{t}\sim\mathcal{D}}\bigg{[}\left(\frac{\tilde{\pi}_{\tilde{\theta}}(a_{t}|s_{t})-\pi_{\theta}(a_{t}|s_{t})}{\pi_{t}(a_{t}|s_{t})}\right)\!\mathop{\min}_{\begin{subarray}{c}x\in 1,2\end{subarray}}Q_{\omega_{x}}({s}_{t},{a}_{t})$
$\qquad+\alpha[H(\tilde{\pi}_{\tilde{\theta}}(\cdot|s_{t}))-H({\pi}_{{\theta}}(\cdot|s_{t}))]\bigg{]}$
Proof.
Following the definition of policy advantage in (14)
$\mathbb{A}^{+}_{\pi}(\tilde{\pi})$
$=$ $\mathop{\mathbb{E}}_{\begin{subarray}{c}s\sim d_{\pi},a\sim\tilde{\pi}\end{subarray}}\left[A_{\pi}(s,a)+\alpha H(\tilde{\pi}(\cdot|s))\right]$
=$\stackrel{{\scriptstyle\mathrm{(a)}}}{{=}}$ $\mathop{\mathbb{E}}_{\begin{subarray}{c}s\sim d_{\pi}\end{subarray}}\left[\sum_{a}\tilde{\pi}(a|s)Q^{\pi}(s,a)-V^{\pi}(s)+\alpha H(\tilde{\pi}(\cdot|s))\right]$
=$\stackrel{{\scriptstyle\mathrm{(b)}}}{{=}}$ $\mathop{\mathbb{E}}_{\begin{subarray}{c}s\sim d_{\pi}\end{subarray}}\Big{[}\sum_{a}\tilde{\pi}(a|s)Q^{\pi}(s,a)-\sum_{a}{\pi}(a|s)Q^{\pi}(s,a)$
$\quad-\alpha H({\pi}(\cdot|s))+\alpha H(\tilde{\pi}(\cdot|s))\Big{]}$
where the equality $(a)$ is according to the state-action advantage value in (2) while $(b)$ follows from (4). Finally, by using Monte Carlo and importance sampling, we can get the lemma.